solve for X: log4 (x^2+3x)=1
Answer by lwsshak3(11628)

solve for X: log4 (x^2+3x)=1
**
log4 (x^2+3x)=1
Convert to exponential form: base(4) raised to log of number(1)=number(x^2+3x)
4^1=(x^2+3x)
x^2+3x-4=0
(x+4)(x-1)=0
x=-4
or
x=1